Find the center and radius of the circle.
64x2 + 64y2 + 16x + 128y + 1 = 0

Answer by lwsshak3(11628)

Find the center and radius of the circle
64x2 + 64y2 + 16x + 128y + 1 = 0
Complete the squares
64(x^2+x/4+1/64)+64(y^2+2y+1)=-1+1+64=64
64(x+1/8)^2+64(y+1)^2=64
(x+1/8)^2+(y+1)^2=1
This is an equation of a circle of standard form: (x-h)^2+(y-k)^2=r^2, (h,k) being the coordinates of the center and r=radius
Center: (-1/8, -1)
Radius=1